OpenStudy (anonymous):

Solve: 4(x + 1) = −4

OpenStudy (anonymous):

4x+4=-4 4x=-8 x=-2

Parth (parthkohli):

A good start is dividing 4 from both sides. \( \color{Black}{\Rightarrow x + 1 = -4 \div 4 }\)

Parth (parthkohli):

\( \color{Black}{\Rightarrow x + 1 = -1}\) right?

mathslover (mathslover):

\[\huge{a(b+c)=ab+ac}\] \[\huge{4(x+1)=-4}\] \[\huge{4x+4=-4}\] \[\huge{4x=-8}\] \[\huge{x=\frac{-8}{4}}\] can u solve this ?

Parth (parthkohli):

@Yashhere can you please not answer the questions? We are an open community for learning. Thanks.

OpenStudy (anonymous):

why are u saying that i have my own space please